triangle A,B,C
side A=376"
side B=388"
side C=? what is side c equal to and what is the equation for that?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

IF THE TEACHER HASN'T EXPLAINED HOW TO SOLVE GEOMETRY PROBLEMS OR YOU DON'T HAVE A TEXT BOOK I SUGGEST YOU GET A BOOK FROM THE SCHOOL OR PUBLIC LIBRARY & DO SOME ADDITIONAL STUDYING.
THE FORMULA FOR THIS PROBLEM IS CALLED THE PYTHAGOREAN THEORUM OR A^2+B^2=C^2
WHERE A & B ARE TWO SIDES & C IS THE HYPOTENUSE.
376^2+388^2=C^2
141,376+150,544=C^2
C^2=291920
C=SQRT291920
C=540.296" THIS IS THE LENGTH OF THE HYPOTENUSE OR THE (C) SIDE.
-----------------------------------------------------------------------
I'M SORRY IT IS LATE & I AM NOT PUNCHING IN THE CORRECT NUMBERS ON MY CALCULATOR.

Answer by stanbon(75887)

Is it a right triangle?
Is C the hypotenuse?
-------------
If C is the hypotenuse:
C= sqrt[376^2 + 388^2]
C = sqrt [291920]
C = 540.296...
Cheers,
Stan H.